# SWBT: Similarity Weighted Behavior Transformer with the Imperfect Demonstration for Robotic Manipulation

### v1

@@ Abstract
Imitation learning (IL), aiming to learn optimal control policies from expert demonstrations, has been an effective method for robot manipulation tasks. However, previous IL methods either only use expensive expert demonstrations and omit imperfect demonstrations or rely on interacting with the environment and learning from online experiences. In the context of robotic manipulation, we aim to conquer the above two challenges and propose a novel framework named Similarity Weighted Behavior Transformer (SWBT). SWBT effectively learn from both expert and imperfect demonstrations without interaction with environments. We reveal that the easy-to-get imperfect demonstrations, such as forward and inverse dynamics, significantly enhance the network by learning fruitful information. To the best of our knowledge, we are the first to attempt to integrate imperfect demonstrations into the offline imitation learning setting for robot manipulation tasks. Extensive experiments on the ManiSkill2 benchmark built on the high-fidelity Sapien simulator and real-world robotic manipulation tasks demonstrated that the proposed method can extract better features and improve the success rates for all tasks. Our code will be released upon acceptance of the paper.
@@ I Introduction
Reinforcement learning (RL) and Imitation Learning (IL) techniques have been extensively researched and applied in the context of robotic manipulation tasks [1, 2, 3, 4, 5, 6] in recent years. RL requires an explicit reward function elaborately pre-defined to guide the agent’s action. Nonetheless, many real-world tasks are intrinsically complicated, making the design of an effective reward function challenging [7, 8]. Imitation learning (IL) aims to directly learn from the demonstrations without designing an explicit reward function and thus is more practical to be applied to real-world applications [9, 10, 11]. Many works of IL require online interactions, which are costly or dangerous, such as robotic manipulation and autonomous driving. On the other hand, offline IL enables data collection in advance, thus independent of online training. Meanwhile, early works of IL normally assume the demonstration is optimal, such as behavioral cloning [12]. The assumption of ideal demonstrations is infeasible or extensively costly since collecting high-quality demonstrations by human labor sometimes would make mistakes. For instance, SayCan [13] collected 276k episodes of data and only retained 12k successful episodes after applying stringent filtering criteria, resulting in more than 90% of the data being wasted.
To address this challenge, recent research in offline IL [14, 15, 16] has emerged, aiming to utilize imperfect demonstrations for agent training. For instance, 2IWIL [14] introduced confidence scores that assess the likelihood of a trajectory being optimal. However, these scores still require manual labeling by human experts for each transition in the imperfect demonstration. DemoDICE [15] formulates an offline IL objective to utilize imperfect demonstrations and mitigate distribution shift from the expert demonstration distribution. However, in the context of robotic manipulation, where data dimensionality is significantly higher, DemoDICE may struggle to accurately align with the data distribution. Another approach, DWBC [17], introduces an additional discriminator to distinguish expert and non-expert demonstrations to mitigate the reward learning as well as leverage imperfect demonstrations. Nonetheless, training the discriminator necessitates costly hyperparameter tuning and struggle with high-dimensional data. Consequently, there is a need for effective offline IL methods capable of leveraging imperfect demonstrations without manual labeling in the context of robotic manipulation with high-dimensional input.
Moreover, recent works [5, 18, 19] have demonstrated that transformers are a suitable option for handling sequential decision-making problems. Recent work Decision Transformer [20] seeks to unify ideas in language modeling and RL by employing the GPT architecture to autoregressively model trajectories and output actions directly. In the context of robotic manipulation settings, we integrate the potent transformer architecture to extract the features of sequential visual-based frames within our proposed framework.
@@
To this end, we propose a novel offline IL-based framework for robotic manipulation, which leveraging imperfect demonstrations, named Similarity Weighted Behavior Transformer (SWBT). Specifically, SWBT contains three main steps. In step 1, we pre-train a multi-modality transformer network using the mixed demonstrations in a self-supervised manner to extract representative features. Since the imperfect demonstrations are readily collected, the mixed demonstrations contain a large amount of fruitful imperfect demonstrations, enabling the transformer to learn large action space knowledge. Inspired by pretraning in Natural Language Processing such as BERT [21], we propose three self-supervised approaches to enhance the pre-trained transformer, including 1) Masked Transition Prediction, 2) Transition Reconstruction, and 3) Action Autoregression. These approaches enable the transformer to take on different roles like a forward dynamic model, inverse dynamic model, and behavior cloning model and thus enhance its ability to extract effective features. In step 2, we aim to identify the useful imperfect demonstrations for training the transformer. Thus, we leverage the pre-trained transformer in step 1 to extract the features and then generate the quality scores by calculating the similarity between the extracted features of the imperfect and expert demonstrations for each state-action pair in the imperfect demonstrations. By doing so, no extra modules, such as discriminator in  [17] are introduced, and the quality scores indicate the quality of the imperfect demonstrations. The simplicity also makes the performance more robust since it does not require much hyper-tuning. In step 3, we perform end-to-end fine-tuning process on the transformer using both expert and imperfect demonstrations. To differentiate the quality of imperfect demonstrations and ensure their beneficial contribution to the transformer, we propose the quality scores calculated in step 2 as weights for the behavior cloning loss. Our contributions as follows.
  • We introduce the Similarity Weighted Behavior Transformer (SWBT). To the best of our knowledge, SWBT is the first work to simultaneously utilize both expert and imperfect demonstrations for training robot manipulation policies within offline imitation learning settings.
  • The robust transformer architecture is leveraged to pre-train an effective feature extractor and introduced quality scores to identify beneficial imperfect demonstrations and eventually improve the success rates of robotic manipulation tasks.
@@
  • Extensive experiments on the ManiSkill2 benchmark demonstrated that SWBT can effectively utilize imperfect demonstrations to improve performance. In addition, our real-world experiments verified that SWBT can be successfully applied to real-world applications.
@@ II Related Work
II-A Offline Imitation Learning with Imperfect Demonstrations
Offline IL has enormous potential to meet critical safety and cost desiderata for real-world applications. Earlier IL method Behavior Cloning (BC) [12] can be used in the offline setting seamlessly and gain comparable performance compared to GAIL methods [22]. IBC [23] and Diffusion Policy [6] further enhance BC by introducing the energy-based models and diffusion models [24]. Another category method, offline inverse reinforcement learning (Offline IRL), focuses on training a discriminator in an adversarial manner [25, 26, 27, 28] or learning a reward function [29] to match the distribution of the expert demonstrations. However, most of these works are based on the assumption that all demonstrations are optimal, thus dropping their performance with imperfect demonstrations.
To mitigate the problem induced by imperfect demonstrations, many imitation learning from imperfect demonstrations methods [14, 30, 31, 32, 29, 33, 34, 35] are proposed. 2IWIL [14] leverages the confidence score to weight the imperfect demonstrations but requires extra human expert annotations. More recently, several works [15, 16, 17] have been proposed to solve the more challenging problem of offline imitation learning with imperfect demonstrations. DWBC [17] combines the above two algorithms by training a discriminator to learn the weights for BC objectives.
II-B Transformers for Robotic Manipulation
Transformers have become the prevalent architecture in various domains. Starting from NLP [36, 37], transformers architectures recently successfully enter the field of vision [38, 39, 40] and achieve impressive results in robot learning [10, 41, 42, 43, 44, 45].  [20, 43, 44, 45] adopt the characteristics of the structure of the transformer to use multi-modality data, such as images and robot states, for policy improvement. Therefore, we leverage the sequential modeling ability of the transformer to improve the performance of the robotic manipulation tasks.
@@ III Methodology
Figure 1: Overview of SWBT. SWBT contains three steps: 1) Transformer Pretraining via Self-Supervised Learning, 2) Calculation of Quality Score by Similarity Metric, and 3) End-to-End Weighted Behavior Cloning Training. We provide an example of the StackCube task, which aims to stack the red cube on the green one. In step 1, we use both expert and imperfect demonstrations to pretrain the transformer. In step 2, we calculate the quality scores for all imperfect segments (e.g., segment $\tau^{imp}_{1}$ and $\tau^{imp}_{2}$ have quality scores of 0.23 and 0.96 respectively. Then, in step 3, the imperfect segment $\tau^{imp}_{1}$ is discarded and the imperfect segment $\tau^{imp}_{2}$ is used for weighted behavior cloning.
@@ III-A Preliminary on Imitation Learning
We consider a standard fully observed Markov decision process (MDP) [46] to model the environment. The MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},r,\rho_{0},\gamma)$ contains the state space $\mathcal{S}$, action space $\mathcal{A}$, reward function $r:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$, state transition function $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$, the initial state distribution $\rho_{0}(s)$, and discount factor $\gamma\in(0,1)$. In our offline IL setting with imperfect demonstrations, the reward function $r$ is unavailable. Instead, we have a static expert demonstration dataset $\mathcal{D}_{e}=\{(s,a,s^{\prime})\}$ and a static imperfect demonstration dataset $\mathcal{D}_{i}=\{(s,a,s^{\prime})\}$. Our goal is to maximize the success rates for the tasks using both the expert dataset and imperfect dataset $\mathcal{D}_{u}=\mathcal{D}_{e}\cup\mathcal{D}_{i}$ without any online interaction with the environments.
III-B Similarity Weighted Behavior Transformer
The overview of the proposed method Similarity Weighted Behavior Transformer (SWBT) is depicted in Figure 1. Specifically, SWBT comprises three main steps. Section III-C details step 1: pre-training a multi-modality transformer network using mixed demonstrations in a self-supervised fashion. Section III-D covers step 2: calculating quality scores based on the similarity between the output features of imperfect and expert demonstrations. Lastly, Section III-E presents step 3: the end-to-end fine-tuning process of the transformer network using weighted behavior cloning.
@@ III-D Calculation of Quality Score by Similarity Metric
To better leverage the high-quality part in the imperfect demonstrations, how to filter them out is a crucial problem. Since the expert demonstrations have high quality, we believe that the imperfect demonstrations that is more similar to expert demonstrations is of higher quality. After pre-training the transformer network in a self-supervised manner, the transformer network now has an enhanced ability to extract expressive and representative features $z^{o}_{i},z^{m}_{i},z^{a}_{i}$ for the input elements. Thus we choose to use the output features of the expert demonstrations and the imperfect demonstrations to calculate the similarities and final quality scores. More specifically, for each trajectory segment $\tau^{imp}$ in the imperfect demonstrations, we extract the output features $z^{o,imp}_{i+l},z^{m,imp}_{i+l},z^{a,imp}_{i+l}$ at the last time step. We also extract the output features $z^{o,exp}_{i+l},z^{m,exp}_{i+l},z^{a,exp}_{i+l}$ of each expert trajectory segment $\tau^{exp}$ in the same way. The similarity between the imperfect features and the expert features are defined as the negative $L_{2}$ distance:
@@
III-E End-to-End Weighted Behavior Cloning Training
After calculating the quality scores for all imperfect trajectory segments $\tau^{imp}$, we can rank them and filter the high-quality part out for learning the optimal policy. In order to prevent low-quality imperfect demonstrations from interfering with policy learning, we define a threshold $\beta$ and choose the imperfect segments $\tau^{imp}$ as follows:
@@ III-E End-to-End Weighted Behavior Cloning Training
For the reserved high-quality segments $\mathcal{D}_{f}$, we use their quality scores as the weights for behavior cloning. Combined with the expert demonstrations $\mathcal{D}_{e}$, the final objective for fine-tuning is given by:
$\mathcal{L}_{\mathrm{full}}=\mathop{{\rm max}}\limits_{\theta}\mathbb{E}_{\tau\sim{\mathcal{D}_{e}}}[logP_{\theta}(a_{i+l}|\tau)]$ (9)
$+\lambda\mathbb{E}_{\tau\sim{\mathcal{D}_{f}}}[q(\tau)\cdot logP_{\theta}(a_{i+l}|\tau)],$
where $\lambda$ is a hyperparameter to balance the training of the expert demonstrations and imperfect demonstrations. We can count the distribution of quality scores $q(\tau)$ offline and then adjust the value of $\beta$ accordingly. We set $l=6,\lambda=0.1,\beta=0.9$ in our implementation.
@@ IV Experiments
We systematically evaluate the Similarity Weighted Behavior Transformer (SWBT) in both simulated and real-world environments. For simulated environments, we built experiments on 5 different tasks on the ManiSkill2 benchmark [47] varying from manipulating rigid and fluid objects, and provided a comprehensive analysis for each component of SWBT with an extensive ablation study. To verify the effectiveness of our algorithm in real-world applications, we built a digital twin system [48] and followed the same setting in the ManiSkill2 benchmark for real-world experiments.
[…]
Figure 3: We conducted experiments on five tasks on the ManiSkill2 benchmark including three rigid-body tasks and two soft-body tasks. The rigid-body tasks include a) PickCube, b) StackCube, and c) PickYCB. The soft-body tasks include d) Fill and e) Hang. The green circles in PickCube and PickYCB represent the goal position.
IV-A Simulation Environments and Datasets
Our simulation experiments are based on the ManiSkill2 benchmark [47], which is built on the high-fidelity Sapien simulator [49] and allows agents to be trained using static demonstrations. The tasks include three rigid object manipulation tasks and two soft-body tasks. For data collection, we trained different level behavior agents using the online DAPG+PPO [50, 51] method provided in the ManiSkill2.

### v2

# Learning from Imperfect Demonstrations with Self-Supervision for Robotic Manipulation

@@ Abstract
Improving data utilization, especially for imperfect data from task failures, is crucial for robotic manipulation due to the challenging, time-consuming, and expensive data collection process in the real world. Current imitation learning (IL) typically discards imperfect data, focusing solely on successful expert data. While reinforcement learning (RL) can learn from explorations and failures, the sim2real gap and its reliance on dense reward and online exploration make it difficult to apply effectively in real-world scenarios. In this work, we aim to conquer the challenge of leveraging imperfect data without the need for reward information to improve the model performance for robotic manipulation in an offline manner. Specifically, we introduce a Self-Supervised Data Filtering framework (SSDF) that combines expert and imperfect data to compute quality scores for failed trajectory segments. High-quality segments from the failed data are used to expand the training dataset. Then, the enhanced dataset can be used with any downstream policy learning method for robotic manipulation tasks. Extensive experiments on the ManiSkill2 benchmark built on the high-fidelity Sapien simulator and real-world robotic manipulation tasks using the Franka robot arm demonstrated that the SSDF can accurately expand the training dataset with high-quality imperfect data and improve the success rates for all robotic manipulation tasks.
@@ I Introduction
Unlike natural language processing (NLP) and computer vision (CV) tasks, where large volumes of text and image data are easily obtained from the web [1, 2], data for robotic manipulation is scarce and precious due to the high-quality data collection process, especially in real-world applications, is extremely challenging, time-consuming, and costly [3]. This makes improving data utilization, particularly incorporating the imperfect data from failed tasks into the model learning process, a critical issue.
@@
As the two dominant learning paradigms, reinforcement learning (RL) and imitation learning (IL) techniques have been extensively researched and applied in the context of robotic manipulation tasks [3, 4, 5, 6, 7, 8] in recent years. However, they still face significant limitations when it comes to utilizing imperfect data. RL requires an explicit reward function elaborately pre-defined to guide the agent’s action. Nonetheless, many real-world tasks are intrinsically complicated, making the design of an effective reward function challenging [9, 10]. Furthermore, RL relies heavily on simulators to facilitate low-cost exploration and learning in a trial-and-error manner, which is impractical in real-world settings due to high labor costs, safety concerns, and physical constraints. On the other hand, IL circumvents the need for explicit reward functions and online exploration by training policy on pre-collected datasets. This makes IL more feasible for real-world applications [11, 12, 13]. However, IL methods, such as behavioral cloning (BC) [14], operate under the assumption that the data used for training is optimal, which is rarely achievable in practice. Various factors, such as physical fatigue, task complexity, and line-of-sight occlusion, can lead to task failures during data collection. For instance, SayCan [15] collected 276k episodes of data and only retained 12k successful episodes after applying stringent filtering criteria, resulting in more than 90% of the data being wasted. In this work, we aim to address the challenge: When the reward function and online exploration are unavailable, how can we leverage imperfect data from task failures to improve the data utilization and success rate for robotic manipulation?
An ideal solution would be to infer a reward function from expert datasets and accurately label all imperfect data. Some offline reinforcement learning algorithms [16, 17, 18, 19, 20] follow this approach and have achieved promising results in state-based environments. However, inferring a reward function without online interaction is challenging. Additionally, most current reward learning algorithms rely on adversarial learning, which is sensitive to hyperparameters, unstable during training, and struggles to learn effectively in high-dimensional environments. Other IL approaches [21, 22, 23, 24, 25, 26] focus on learning a policy that mimics the expert’s behavior, aiming to mitigate the distribution shift problem [27] and improve model performance. However, in the context of robotic manipulation, where expert data are scarce and data dimensionality is significantly higher, many methods struggle to align accurately with the data distribution. For instance, training the discriminator in DWBC [24] necessitates costly hyperparameter tuning and struggles with high-dimensional data. Consequently, there is a need for effective offline IL methods capable of leveraging imperfect demonstrations with high-dimensional input.
Our key insight is motivated by the observation: Many steps in a failed trajectory are of high quality despite the overall task failure. For example, when attempting to open a cabinet, the robotic arm may successfully grasp the handle but fail to open the cabinet fully. In such cases, the failed data leading up to the successful grasp is valuable and should be regarded as successful data. More importantly, the expert data from task success certainly contains data that succeeded in grasping the handler, so these two pieces of data will be very similar, providing hints as to how to select high-quality imperfect data.
@@
To this end, we propose a novel Self-Supervised Data Filtering framework (SSDF) for robotic manipulation, which extends the training dataset using high-quality imperfect data filtered based on the similarity between the failed trajectory segment and the expert dataset. Specifically, SSDF contains three main steps. In step 1, we pre-train a multi-modality transformer network using the mixed demonstrations in a self-supervised manner to extract representative features. Inspired by pertaining process in BERT [28], we propose three self-supervised objectives to enhance the pre-trained transformer, including 1) Masked Transition Prediction, 2) Transition Reconstruction, and 3) Action Autoregression. These approaches enable the transformer to take on different roles like a forward dynamic model, inverse dynamic model, and behavior cloning model and thus enhance its ability to extract effective features. In step 2, we aim to identify the useful imperfect demonstrations according to the quality scores and use them to augment the original expert dataset. Thus, we leverage the pre-trained transformer in step 1 to extract the features and then generate the quality scores by calculating the similarity between the extracted features of the imperfect and expert demonstrations for each state-action pair in the imperfect demonstrations. In step 3, we perform weighted behavior cloning for the high-quality imperfect data, and the original expert dataset can be used for any IL method. To differentiate the quality of imperfect demonstrations and ensure their beneficial contribution to the downstream policy, we propose the quality scores calculated in step 2 as weights for the behavior cloning loss. Our contributions are as follows.
  • We introduce the Self-Supervised Data Filtering framework (SSDF), which can utilize expert and imperfect demonstrations to train robot manipulation policies.
  • We design three self-supervised objectives for feature extraction and a novel similarity calculation method, which can accurately expand the training dataset with high-quality imperfect data.
@@
  • Extensive experiments on the ManiSkill2 and real-world tasks demonstrated that SSDF can effectively utilize imperfect demonstrations to improve performance.
@@ II Related Work
Imitation Learning has enormous potential to meet critical safety and cost desiderata for real-world applications. As an important class of methods, Behavior Cloning (BC) [14] can be used in the offline setting seamlessly and gain comparable performance compared to GAIL methods [29]. IBC [30], ACT [31], and Diffusion Policy [8] further enhance BC using the network with better expressive capability and advanced training objectives. However, most of these works are based on the assumption that all demonstrations are optimal, thus dropping their performance with imperfect demonstrations.
Some offline reinforcement learning algorithms [16, 17, 18, 19, 20] follow this approach and have achieved promising results in state-based environments. However, inferring a reward function without online interaction is challenging. Additionally, most current reward learning algorithms rely on adversarial learning, which is sensitive to hyperparameters, unstable during training, and struggles to learn effectively in high-dimensional environments.
Offline inverse reinforcement learning (Offline IRL) focuses on training a discriminator in an adversarial manner [32, 33, 34, 35] or learning a reward function [16, 18, 19, 20] to match the distribution of the expert demonstrations. For example, MIIRL [36] proposes to train a conservative world model for learning the reward function from the expert data. Nonetheless, the adversarial optimization process in these methods can lead to training instability and collapse, which is exacerbated in high-dimensional environments. In this work, we aim to conquer the challenge of leveraging imperfect data in the context of robotic manipulation with high-dimensional inputs.
Many IL methods with imperfect demonstrations [21, 37, 38, 39, 40, 41, 42] are also proposed. Nonetheless, most methods require extra labeling or online interaction to provide more information to learn. More recently, several works [23, 25, 24, 26] have been proposed to solve the more challenging problem of offline imitation learning with imperfect demonstrations. ISW-BC [26] proposes an importance-sampling-based method to re-weigh imperfect data and implicitly make the learned policy close to the expert policy. Some other imitation learning methods leveraging powerful architectures [43, 44], such as transformers [45], can flexibly integrate diverse data. These methods typically use goal-conditioned training [46, 47] or pretraining [48, 49, 50] to extract more representative features. However, they are not specifically designed to handle imperfect data. In contrast, our method designs a tailored self-supervised framework for computing the quality scores of imperfect data and is applied to vision-based robotic manipulation.
@@ III Methodology
Figure 1: Overview of SSDF. SSDF contains three steps: 1) Transformer Pretraining via Self-Supervised Learning, 2) Calculation of Quality Score by Similarity Metric, and 3) Policy Learning with Weighted Behavior Cloning.
@@ III-A Preliminary on Imitation Learning
We consider a standard fully observed Markov decision process (MDP) [51] to model the environment. The MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},r,\rho_{0},\gamma)$ contains the state space $\mathcal{S}$, action space $\mathcal{A}$, reward function $r:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$, state transition function $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$, the initial state distribution $\rho_{0}(s)$, and discount factor $\gamma\in(0,1)$. In the IL setting with imperfect demonstrations, the reward function $r$ is unavailable. Instead, we have a static expert demonstration dataset $\mathcal{D}_{e}=\{(s,a,s^{\prime})\}$ and a static imperfect demonstration dataset $\mathcal{D}_{i}=\{(s,a,s^{\prime})\}$. Our goal is to maximize the success rates for the tasks using both the expert dataset and imperfect dataset $\mathcal{D}_{u}=\mathcal{D}_{e}\cup\mathcal{D}_{i}$ without any reward information and online interaction with the environments.
III-B Self-Supervised Data Filtering Framework
The overview of the proposed method Self-Supervised Data Filtering framework (SSDF) is depicted in Figure 1. Specifically, SSDF comprises three main steps. (1) Transformer Pretraining via Self-Supervised Learning. (2) Calculation of Quality Score by Similarity Metric. (3) Policy Learning with Weighted Behavior Cloning.
@@ III-D Calculation of Quality Score by Similarity Metric
Since the expert demonstrations have high quality, we believe that the imperfect demonstrations that is more similar to expert demonstrations is of higher quality. After pre-training the transformer network in a self-supervised manner, the transformer network now has an enhanced ability to extract expressive and representative features $z^{o}_{i},z^{m}_{i},z^{a}_{i}$ for the input elements. Thus, we choose to use the output features of the expert demonstrations and the imperfect demonstrations to calculate the similarities and final quality scores. More specifically, for each trajectory segment $\tau^{imp}$ in the imperfect demonstrations, we extract the output features $z^{o,imp}_{i+l},z^{m,imp}_{i+l},z^{a,imp}_{i+l}$ at the last time step. We also extract the output features $z^{o,exp}_{i+l},z^{m,exp}_{i+l},z^{a,exp}_{i+l}$ of each expert trajectory segment $\tau^{exp}$ in the same way. The similarity between the imperfect features and the expert features are defined as the negative $L_{2}$ distance:
@@
III-E Policy Learning with Weighted Behavior Cloning
After calculating the quality scores for all imperfect trajectory segments $\tau^{imp}$, we can rank them and filter the high-quality part out to learn the optimal policy. In order to prevent low-quality imperfect demonstrations from interfering with policy learning, we define a threshold $\beta$ and choose the imperfect segments $\tau^{imp}$ as follows:
@@ III-E Policy Learning with Weighted Behavior Cloning
For the reserved high-quality dataset $\mathcal{D}_{f}$, we use their quality scores as the weights for behavior cloning. Combined with the expert demonstrations $\mathcal{D}_{e}$, the final objective for downstream policy learning is given by:
$\mathcal{L}_{\mathrm{full}}=\mathop{{\rm min}}\limits_{\theta^{{}^{\prime}}}\mathcal{L}_{\mathrm{policy}}-\lambda\mathbb{E}_{\tau\sim{\mathcal{D}_{f}}}[q(\tau)logP_{\theta^{{}^{\prime}}}(a_{i+l}|\tau)],$ (9)
where $\mathcal{L}_{\mathrm{policy}}$ is the objective for any downstream policy model parameterized by $\theta^{{}^{\prime}}$ on the expert dataset $\mathcal{D}_{e}$, $\lambda$ is a hyperparameter to balance the training of the expert demonstrations and imperfect demonstrations. We finetune the transformer in our simulation experiments and use ACT [31] in our real-world experiments. We can count the distribution of quality scores $q(\tau)$ offline and then adjust the value of $\beta$ accordingly. We set $l=6,\lambda=0.1,\beta=0.9$ in our implementation.
@@ IV Experiments
Figure 3: Experiment Setup: we conducted experiments on five tasks on the ManiSkill2 benchmark and five tasks using the real-world single Franka robotic arm. Fill and Hang are soft-body tasks. The green circles in PickCube and PickYCB represent the goal position.
IV-A Experiment Setup
Simulation Experiments. Our simulation experiments are based on the ManiSkill2 benchmark [52], which is built on the high-fidelity Sapien simulator [53] and allows agents to be trained using static demonstrations. The tasks include three rigid object manipulation tasks and two soft-body tasks as shown in Figure 3. For data collection, we trained different level behavior agents using the online DAPG+PPO [54, 55] method provided in the ManiSkill2.